1.25 x 1022 atoms Cl


1
Expert's answer
2021-03-31T08:00:44-0400

As 1 mol of a substance contains 6.022 × 1023 particles, 1.25 × 1022 atoms of Cl equals:

n = 1.25 × 1022 / 6.022 × 1023 = 2.08 × 10-2 mol


Answer: 2.08 × 10-2 mol